### v1

# SelfCite: Self-Supervised Alignment for Context Attribution in Large Language Models

Abstract
We introduce SelfCite, a novel self-supervised approach that aligns LLMs to generate high-quality, fine-grained, sentence-level citations for the statements in their generated responses. Instead of only relying on costly and labor-intensive annotations, SelfCite leverages a reward signal provided by the LLM itself through context ablation: If a citation is necessary, removing the cited text from the context should prevent the same response; if sufficient, retaining the cited text alone should preserve the same response. This reward can guide the inference-time best-of-N sampling strategy to improve citation quality significantly, as well as be used in preference optimization to directly fine-tune the models for generating better citations. The effectiveness of SelfCite is demonstrated by increasing citation F1 up to 5.3 points on the LongBench-Cite benchmark across five long-form question answering tasks.
[…]
Yung-Sung Chuang , Shang-Wen Li , Scott Yih
1 Introduction
Assistants built using large language models (LLMs) have become ubiquitous in helping users gather information and acquire knowledge (OpenAI, 2022, 2023). For instance, when asked about recent news, an assistant can read through dozens of relevant articles—potentially more than a user could comb through themselves—and use these articles as context to provide a clear, specific answer to the user’s query. While this ability can greatly accelerate information gathering, LLMs often produce hallucinations—content that sounds plausible but is actually fabricated (Ji et al., 2023). Even when provided with accurate context, models may misinterpret the data or include details that are not supported by the context (Shi et al., 2024; Chuang et al., 2024).
Although completely eliminating hallucinations remains difficult, existing approaches have sought to enhance the reliability of LLMs by providing context attributions–commonly referred to as citations–which are fine-grained references to relevant evidences from the context, alongside generated responses for user verification (Menick et al., 2022; Slobodkin et al., 2024; Zhang et al., 2024). While they have shown promise in generating citations, an outstanding challenge is their reliance on annotated data either from human (Menick et al., 2022; Slobodkin et al., 2024) or costly proprietary APIs (Zhang et al., 2024; Huang et al., 2024) to train models to generate citations. Collecting annotations can be time-consuming and costly, especially with long-context documents.
To address this challenge, we introduce SelfCite, a novel alignment approach designed to autonomously enhance the quality of citations generated by LLMs without the need for any annotations in the alignment process. Drawing inspiration from model interpretability techniques (Lei et al., 2016; Cohen-Wang et al., 2024), SelfCite leverages the inherent capabilities of LLMs to provide feedback through context ablation—a process to evaluate the necessity and sufficiency of a citation. If removing the cited text prevents the LLM from assigning high probability to the same response, we can infer that it is necessary for the LLM. Conversely, if the response remains highly probable despite removing all context other than the cited text, this indicates that the citation is sufficient for the LLM to make the claim. This self-evaluation mechanism enables SelfCite to calculate a reward signal without relying on the annotation processes.
Building on this intuition, we design a reward that can be cheaply computed by the LLM itself, composed by probability drop and probability hold in context ablation. By integrating this reward function into a best-of-N sampling strategy, SelfCite achieves substantial improvements in citation quality. Furthermore, we employ this reward for preference optimization using SimPO (Meng et al., 2024), which not only maintains these improvements but also eliminates the need for best-of-N sampling. We outperform the previous state of the art on the LongBench-Cite benchmark (Zhang et al., 2024) by up to 5.3 points in F1 scores, and showing a promising direction to bootstrap the citation quality from LLMs via self-rewarding.
Figure 1: The SelfCite framework calculates rewards based on two metrics: necessity score (probability drop) and sufficiency score (probability hold). First, the full context is used to generate a response. Then, the framework evaluates the probability of generating the same response after (1) removing the cited sentences from the context and (2) using only the cited sentences in the context. The probability drop and hold are computed from these probability differences, and their sum is used as the final reward.
2 Method
In this section, we describe the SelfCite framework. We begin by introducing the task of generating responses with context attributions (2.1), referred to as citations for brevity. We then design a reward for providing feedback on citation quality without human annotations (2.2) as illustrated in Fig. 1. Finally, we discuss two approaches for utilizing this reward to improve citation quality: best-of-N sampling (2.3) and preference optimization (2.4).
2.1 Problem Formulation
We first formalize the task of generating responses with context attributions and the metrics to self-evaluate context attributions within the SelfCite framework, inspired by previous papers (Zhang et al., 2024; Cohen-Wang et al., 2024) but adapted to our proposed self-supervised reward.
Setup.
Consider employing an autoregressive language model (LM) to generate a response to a specific query given a context of relevant information. Specifically, given an LM $p_{\text{LM}}$, let $p_{\text{LM}}(t_{i}\mid t_{1},\ldots,t_{i-1})$ denote its output distribution over the next token $t_{i}$ based on a sequence of preceding tokens $t_{1},\ldots,t_{i-1}$. Next, let $C$ represent the context of relevant information. This context is partitioned into $|C|$ sentences: $c_{1},c_{2},\dots,c_{|C|}$. Each sentence $c_{j}$ is prepended with a unique identifier (e.g. sentence index $j$) as a way for the model to reference the sentence when generating citations. The context $C$ is followed by a query $Q$, a question or instruction for the model. A response $R$ is then sampled from the LM $p_{\text{LM}}$.
Generating Responses with Context Attributions.
In SelfCite, following prior work on generating responses with context attributions (Zhang et al., 2024), each statement $r_{i}$ in the response $R$ is followed by a citation sequence $e_{i}$ consisting of the identifiers of sentences from the context $C$. Thus, the entire response sequence $R$ is $\{r_{1},e_{1},r_{2},e_{2},\dots,r_{S},e_{S}\}$, where $S$ is the total number of generated statements. The citation $e_{i}$ is intended to reference sentences that support the generation of $r_{i}$. Formally, for each response statement $r_{i}$, the model outputs a citation sequence $e_{i}=\{e_{i}^{1},e_{i}^{2},\dots,e_{i}^{m}\}$, where each $e_{i}^{j}\in\{1,2,\dots,|C|\}$ corresponds to a specific sentence number in the context $C$, and $m$ sentences are cited. Note that this citation sequence may be empty. The entire response $R$ consisting of statements $r_{i}$ followed by citations $e_{i}$ is sampled from the LM $p_{\text{LM}}$ as follows:
$r_{i}$ $\sim p_{\text{LM}}\left(\cdot\mid c_{1},\ldots,c_{|C|},Q,r_{1},e_{1},\ldots,r_{i-1},e_{i-1}\right),$
$e_{i}$ $\sim p_{\text{LM}}\left(\cdot\mid c_{1},\ldots,c_{|C|},Q,r_{1},e_{1},\ldots,r_{i-1},e_{i-1},r_{i}\right).$
The objective of optimizing the LM is to ensure that the citation sequence $e_{i}$ accurately reflects the evidence from the context that supports the generation of $r_{i}$. In the SFT setting (Zhang et al., 2024), the probability of a “ground truth” annotated responses and citations $\{\hat{r}_{1},\hat{e}_{1},...,\hat{r}_{S},\hat{e}_{S}\}$ will be maximized, given the input $C$ and $Q$, but it is not trivial to do further alignment with feedback after the SFT data is used up. To achieve this, we introduce SelfCite that can evaluate the quality of these citations based on context ablation as a reward for further preference optimization.
2.2 Self-Supervised Reward via Context Ablation
We measure the quality of a citation sequence $e_{i}$ by the changes in the LM’s probability of generating $r_{i}$ when the cited sentences are either removed from or isolated within the context. To simplify the notation, let all the cited context sentences be $E_{i}=\{c_{e_{i}^{1}},c_{e_{i}^{2}},\dots,c_{e_{i}^{m}}\}$. We define two key metrics: necessity score and sufficiency score, and finally combine them into the final reward, as shown in Fig. 1.
Necessity Score: Probability Drop.
This metric quantifies the decrease in the probability of generating $r_{i}$ when the cited sentences in $E_{i}$ are all removed from the context (denoted as set minus $\setminus$ operator). Formally, it is defined as:
[…]
To keep the equation concise, we ignore $Q$ and $\{r_{1},e_{1},...,r_{i-1},e_{i-1}\}$ in the equation, but they are staying in the context history when computing the probabilities. A larger probability drop indicates that the removal of $E_{i}$ significantly diminishes the likelihood of generating $r_{i}$, thereby validating the necessity of the cited evidence.
Sufficiency Score: Probability Hold.
Conversely, this metric measures if the probability of generating $r_{i}$ is still kept large when only the cited sentences are kept in the context, effectively testing the sufficiency of the citation to support the response statement. Formally:
[…]
A more positive value of probability hold indicates that the cited sentences alone are sufficient to support the generation of $r_{i}$, while removing all the other irrelevant context. Please note that the values of probability drop or hold can be either positive or negative. For example, if the citation is not relevant to $r_{i}$ or even distracting, it is possible for $p(r_{i}\mid E_{i})$ to be lower than $p(r_{i}\mid C)$.
Final Reward.
To comprehensively evaluate the necessity and sufficiency of the generated citations, we add the two metrics together, where the opposing terms cancel out:
$\text{Reward}(e_{i})=\log p_{\text{LM}}\left(r_{i}|E_{i}\right)-\log p_{\text{LM}}\left(r_{i}|C\setminus E_{i}\right).$ (1)
The combined reward measures if the citations are both necessary and sufficient for generating the response $r_{i}$.
2.3 Best-of-N Sampling
To leverage the self-supervised reward computed via context ablation, we employ a best-of-N sampling strategy, which is a common way to test the effectiveness of a reward design (Gao et al., 2023a; Lightman et al., 2024) as a performance oracle without any confounders from training. After generating the full response, we locate the position where the citation tags <cite>...</cite> are generated. Within the citation tags, we sample $N$ candidate citation sequences and select the citation set that maximizes the combined reward metric, Eq. (1). The corresponding procedure is shown in Algorithm 1. After obtaining the selected citations $\{e_{1}^{*},\dots,e_{S}^{*}\}$, we replace the original citation sequence $e_{i}$ with the optimal citation set $e_{i}^{*}$ for the response statement $r_{i}$, while keeping the response statements $\{r_{1},\dots,r_{S}\}$ unchanged. This process is repeated for each statement in the response $R$ to obtain the final, citation-improved output $R^{*}=\{r_{1},e_{1}^{*},\dots,r_{S},e_{S}^{*}\}$. To prevent the model from citating too many sentences, we exclude the BoN candidates that cites more than 384 tokens in total, unless they are all from a single sentence.
Algorithm 1 SelfCite Best-of-N Sampling for Citations
0:  LM $p_{\text{LM}}$, context $C$, query $Q$, response $R$, number of candidates $N$
  for $r_{i}\in R$ do
     for $k=1,\dots,N$ do
        $e_{i}^{(k)}\sim p_{\text{LM}}(\cdot\mid r_{i},C,Q)$
        $\mathrm{Reward}(e_{i}^{(k)})$ $=\log p_{\text{LM}}\bigl{(}r_{i}\mid E_{i}^{(n)}\bigr{)}-\log p_{\text{LM}}\bigl{(}r_{i}\mid C\setminus E_{i}^{(n)}\bigr{)}$
     end for
     $e_{i}^{*}=\arg\max_{k}\;\mathrm{reward}\bigl{(}e_{i}^{(k)}\bigr{)}$
  end for
  return $R^{*}=\{r_{1},e_{1}^{*},\dots,r_{S},e_{S}^{*}\}$
2.4 Preference Optimization
Best-of-N sampling is a straightforward way to obtain better citations, but at the additional inference time of generating candidates and reranking. Thus, we try to internalize the ability of generating better citations back to the LM itself.
Given documents and queries, we can prompt the LM to generate the responses along with the citations $R=\{r_{1},e_{1},...,r_{S},e_{S}\}$. By further applying best-of-N sampling, we can obtain new responses of the same statements but with better citations $R^{*}=\{r_{1},e^{*}_{1},...,r_{S},e^{*}_{S}\}$. Such preference data can be used in direct preference optimization (DPO) (Rafailov et al., 2024) to align the model based on the preference between the original outputs and improved outputs. DPO typically requires more memory usage than SFT due to the need of a reference model. Also, optimizing with preference data pairs inherently makes the per-GPU batch size to be at least 2, limiting the maximum context length that can be used. To address these challenges, we use SimPO (Meng et al., 2024), a variant of DPO that does not require a reference model, freeing up memory for long-context fine-tuning. We further apply Liger-Kernel (Hsu et al., 2024), a collection of efficient Triton kernels, to optimize memory usage and scale the context length to 25.6K, as detailed in Appendix 7. Through this self-supervised alignment process, which does not require ground-truth answers or human annotations, the model learns to generate more accurate and contextually grounded citations on its own.
3 Experiments
We evaluate the effectiveness of SelfCite by applying the best-of-N sampling and preference optimization methods to existing models that generate responses with citations.
3.1 Model Details
We use the Llama-3.1-8B model (Dubey et al., 2024) fine-tuned on LongCite-45K SFT data, namely the LongCite-8B model (Zhang et al., 2024) as the start point for both best-of-N sampling and preference optimization. We adopt the same text segmentation strategy from Zhang et al. (2024): each document is split into individual sentences using NLTK (Bird, 2006) and Chinese punctuations, and each sentence is prepended with a unique identifier in <C{$i$}> format. These identifiers serve as the citation indices, enabling the model to cite relevant context right after the statements with the format of <statement> {content ...} <cite>[$i_{1}-i_{2}$][$i_{3}-i_{4}$]...</cite></statement>. This format allows the model to cite a single sentence (e.g. $i_{1}=i_{2}$) or a span (e.g. $i_{1}<i_{2}$) efficiently within several tokens. The responses are generated via top-p sampling (Holtzman et al., 2020) with p=0.7 and temperature=0.95. We set p=0.9 and temperature=1.2 when doing best-of-N sampling for the citation strings to increase the diversity. We set N=10 in all the experiments considering the limited diversity in citations. [^1] After deduplicating repeated citation candidates, on average there are only 4.8 candidates left per statement in the BoN experiment on LongBench-Cite, with a standard deviation of 3.2.
3.2 Preference Optimization
LongCite-45K.
Best-of-N sampling (Section 2.3) requires no training, so no training data is used. For preference optimization with SimPO (Section 2.4), we use 2K document–question pairs from LongCite-45K (Zhang et al., 2024) as the training set but we do not use its ground-truth responses with high-quality citations for SFT. Instead, we generate model responses from the documents and queries, then apply best-of-N to refine citations. We label the original responses as rejected and replace their citations with BoN-refined ones to create the chosen responses, forming preference pairs to build the dataset for SimPO.
Data Construction and Length Balancing
Since best-of-N responses tend to have slightly longer citations, directly fine-tuning on them can lead the model to adopt a shortcut—generating longer citations instead of improving citation quality. To prevent this, we introduce length balancing: if an original response has a shorter citation length than the best-of-N response, we insert random citations from nearby sentences. This encourages the model to focus on where to cite rather than simply citing more. Details are provided in Appendix 9, with an ablation study in Section 4.2.
3.3 Evaluation
Benchmark.
We evaluate our approach on LongBench-Cite (Zhang et al., 2024), a comprehensive benchmark specifically designed for long-context QA with citations (LQAC). Given a long context $C$ and a query $Q$, the model must produce a multi-statement answer with each statement cites relevant supporting sentences in $C$. Unlike chunk-level citation schemes (Gao et al., 2023b) which cites short paragraphs, LongBench-Cite adopts sentence-level citations to ensure semantic integrity and finer-grained evidence tracking. LongBench-Cite assesses two main aspects:
  • Citation Quality: Whether each statement is fully supported by relevant and only relevant sentences. GPT-4o measures citation recall (extent to which a statement is fully or partially supported by the cited text) and citation precision (whether each cited text truly supports the statement). These are combined into a citation F1 score. Additionally, we track average citation length (tokens per citation) to promote fine-grained citations over unnecessarily long passages.
  • Correctness: How accurately and comprehensively the response answers the query disregarding the citations. This is scored by GPT-4o in a zero-/few-shot fashion based on the query and reference answers.
The benchmark contains five datasets, including single-doc QA MultiFieldQA-en/zh (Bai et al., 2023), multi-doc QA HotpotQA (Yang et al., 2018) and DuReader (He et al., 2018), one summarization dataset GovReport (Huang et al., 2021), and LongBench-Chat (Bai et al., 2024) which covers diverse real-world queries with long contexts such as document QA, summarization, and coding.
Baselines.
SelfCite is compared with these baselines.
  • Prompting: Zhang et al. (2024) propose the baseline of prompting LLMs with an one-shot example. This can be applied to proprietary models including GPT-4o (OpenAI, 2023), Claude-3-sonnet (Anthropic, 2024), and GLM-4 (GLM et al., 2024), as well as open-source models including GLM-4-9B-chat (GLM et al., 2024), Llama-3.1-{8,70}B-Instruct (Dubey et al., 2024), and Mistral-Large-Instruct (Mistral, 2024).
  • Contributive context attribution: Contributive context attribution seeks to directly identify the parts of the context that cause the model to generate a particular statement. We consider ContextCite (Cohen-Wang et al., 2024), a contributive context attribution method that performs several random context ablations to model the effect of ablating different parts of the context on a generated statement. We apply ContextCite with 32 and 256 times of random context ablations, with the details described in Appendix 8.
  • Fine-tuned models: LongCite-8B and 9B released by Zhang et al. (2024), trained on LongCite-45K, fine-tuned from Llama-3.1-8B (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024), respectively.
Table 1: Citation recall (R), citation precision (P), citation F1 (F1), and citation length evaluated on LongBench-Cite benchmark. The best of our results are bolded. The best of previous state of the art are underlined. indicates the results taken from Zhang et al. (2024). Our repro. means our reproduced results.
Model Longbench-Chat MultifieldQA HotpotQA Dureader GovReport Avg. Citation
R P F1 R P F1 R P F1 R P F1 R P F1 F1 Length
Proprietary models
GPT-4o 46.7 53.5 46.7 79.0 87.9 80.6 55.7 62.3 53.4 65.6 74.2 67.4 73.4 90.4 79.8 65.6 220
Claude-3-sonnet 52.0 67.8 55.1 64.7 85.8 71.3 46.4 65.8 49.9 67.7 89.2 75.5 77.4 93.9 84.1 67.2 132
GLM-4 47.6 53.9 47.1 72.3 80.1 73.6 47.0 50.1 44.4 73.4 82.3 75.0 82.8 93.4 87.1 65.4 169
Open-source models
GLM-4-9B-chat 25.9 20.5 16.7 51.1 60.6 52.0 22.9 28.8 20.1 45.4 48.3 40.9 5.7 8.2 6.3 27.2 96
Llama-3.1-8B-Instruct 14.1 19.5 12.4 29.8 44.3 31.6 20.2 30.9 20.9 22.0 25.1 17.0 16.2 25.3 16.8 19.7 100
Llama-3.1-70B-Instruct 25.8 32.0 23.2 53.2 65.2 53.9 29.6 37.3 28.6 38.2 46.0 35.4 53.4 77.5 60.7 40.4 174
Mistral-Large-Instruct 19.8 23.9 19.0 71.8 80.7 73.8 34.5 40.9 32.1 58.3 67.0 60.1 67.9 79.6 72.5 51.5 132
Contributive context attribution (with Llama-3.1-8B-Instruct)
ContextCite (32 calls) 56.7 76.8 58.0 76.1 87.2 78.9 40.5 54.7 43.9 58.0 82.4 65.0 67.1 88.8 75.6 64.3 92.7
ContextCite (256 calls) 63.5 83.1 64.7 78.8 89.8 81.8 46.5 60.8 49.2 61.7 89.1 70.1 69.1 93.5 78.8 68.9 100.8
Fine-tuned models
LongCite-9B 57.6 78.1 63.6 67.3 91.0 74.8 61.8 78.8 64.8 67.6 89.2 74.4 63.4 76.5 68.2 69.2 91
LongCite-8B 62.0 79.7 67.4 74.7 93.0 80.8 59.2 72.1 60.3 68.3 85.6 73.1 74.0 86.6 78.5 72.0 85
Ours: SelfCite
LongCite-8B (Our repro.) 67.0 78.1 66.6 74.8 90.7 79.9 60.8 77.9 64.1 67.1 87.2 73.7 81.6 89.3 84.5 73.8 83.5
+ BoN 68.4 81.3 71.2 76.1 92.8 81.2 67.2 81.0 68.8 70.6 90.9 76.9 87.6 92.4 89.3 77.5 93.4
+ SimPO 68.1 79.5 69.1 75.5 92.6 81.0 69.4 82.3 71.5 72.7 91.6 78.9 86.4 92.9 89.1 77.9 105.7
+ SimPO then BoN 73.3 79.4 72.8 76.7 93.2 82.2 69.4 83.0 71.1 74.2 92.2 80.3 86.7 92.7 89.2 79.1 94.7
Llama-3.1-8B-Instruct (fully self-supervised setting)
+ SFT on ContextCite 52.3 70.6 56.5 79.1 90.5 82.0 54.5 72.3 56.3 54.9 79.0 61.6 63.7 84.9 72.3 65.7 83.0
       + BoN 54.8 67.6 58.1 80.4 90.5 83.0 58.3 70.0 57.5 57.6 79.0 63.1 67.2 84.8 74.6 67.3 80.4
       + SimPO 63.3 74.3 64.6 80.2 88.9 82.4 59.7 76.9 61.0 59.0 80.9 65.4 68.5 86.6 76.1 69.9 90.2
       + SimPO then BoN 66.0 82.4 71.1 81.5 90.7 83.2 61.3 70.0 59.9 62.1 81.4 67.4 68.8 86.2 76.1 71.5 87.4
3.4 Main Results
Citation Quality.
Table 1 presents our main results. Our best-of-N sampling (BoN) consistently improves both citation recall and citation precision across tasks, increasing the overall F1 score from 73.8 to 77.5. Using SimPO to internalize BoN’s gains—eliminating the need for BoN sampling—achieves a similar improvement, with an F1 of 77.9. Applying BoN again to the SimPO fine-tuned model further boosts F1 by 5.3 points to 79.1, the highest across the datasets, suggesting room for further gains. Our results surpass LongCite-8B/9B at similar citation lengths and outperform proprietary model prompting while producing shorter citations. Compared to ContextCite, which also relies on context ablation, our approach is significantly better. A key reason is that ContextCite estimates sentence importance from scratch using linear regression, while we rerank existing LLM-generated citation candidates, leading to more efficient and accurate citation quality estimation. Additionally, we evaluate the latest released Claude Citations API, as shown in Appendix 10 that SelfCite achieves strong results very close to this commercial-level API that specialized for providing citations.
Fully Self-Supervised Setting.
In our main experiment, we start from the Llama-3.1-8B model fine-tuned on the LongCite-45K SFT data, which effectively kick-starts its ability to generate structured citations for best-of-N sampling. The subsequent SimPO alignment stage is entirely self-supervised. We are also curious if it is possible to start from a fully self-supervised SFT model and then apply our self-supervised alignment after that. To begin with, we automatically generate 11K citation SFT data using ContextCite (see Appendix 8 for details) to replace the LongCite-45K annotations in the training data, as shown in the results at the bottom of Table 1. We can see that SFT on ContextCite can achieve decent initial results (65.7 F1) but still far from LongCite-8B (73.8 F1). BoN helps improving F1 to 67.3. After SimPO training, it achieves 69.9 F1, and additionally applying BoN can boost its F1 by 5.8 to 71.5, significantly closing the gap to LongCite-8B, showing our alignment method not only improve the supervised models, but also enhance the models purely trained from self-supervision.
Table 2: Answer correctness when responding with or without citations. indicates results taken from Zhang et al. (2024). The header contains abbreviations for the same five datasets in Table 1.
Model Long. Multi. Hot. Dur. Gov. Avg
Answering without citations
LongSFT-8B 68.6 83.6 69.0 62.3 54.4 67.6
LongSFT-9B 64.6 83.3 67.5 66.3 46.4 65.6
Llama-3.1-8B-Instruct 61.6 73.3 64.5 39.4 62.1 60.2
Answering with citations
LongCite-8B (Our repro.) 67.6 86.7 69.3 64.0 60.4 69.6
+ SimPO 67.4 86.7 67.5 66.0 61.3 69.8
Llama-3.1-8B-Instruct 67.4 87.9 73.5 67.8 62.1 71.7
+ SFT on ContextCite 58.8 83.4 65.8 57.8 57.5 64.6
       + SimPO 56.8 80.9 65.3 59.5 60.9 64.7
Answer Correctness.
For best-of-N sampling, only the citation parts are modified, so the responses it generates to answer the questions are the same as those of the original LongCite-8B model, maintaining the same correctness. For the SimPO finetuned models, we test their answer correctness by the evaluation in Zhang et al. (2024), which contains two settings: answering with/without citations. If answering with citations, the citation parts will be removed when evaluating the answer correctness. The results in Table 2 show that the SimPO finetuning does not change the correctness of the LongCite-8B model much. The correctness is similar to LongSFT-8B/9B (Zhang et al., 2024), which are ablation baselines finetuned on LongCite-45k QA pairs but without the citation parts. When finetuning Llama-3.1-8B-Instruct with ContextCite SFT data, the answer correctness slightly degrades. It is probably caused by the fact that we do not mix any instruction following dataset in the SFT stage as Zhang et al. (2024) does. However, the further SimPO step does not change the answer correctness significantly, while lead to better citations for user verification. Since the SFT stage is not our main method here, we leave the better data mixture strategy in SFT stage in the future work.
4 Analysis
4.1 Ablation Study on Rewards
To better understand our final reward design, we explore various reward strategies in the BoN sampling process. Here, all BoN candidates are pre-generated and fixed, the reward is the only factor affecting results. Table 3 presents our ablation results on HotpotQA, while citation lengths are computed across all LongBench-Cite datasets for direct comparison with Table 1. We evaluate four alternative reward designs. BoN by LM log prob re-ranks candidates simply by the probability of the citation string, <cite>[$i_{1}-i_{2}$][$i_{3}-i_{4}$]...</cite>, which is similar to the effects of beam search. We observe that this strategy slightly boosts recall while reducing precision, resulting in a minor reduction in F1. BoN by max citation length always selects the candidates with the longest citations, i.e. citing the greatest number of sentences. Although it improves recall, it significantly reduces precision from 77.9 to 73.6 and inflates the citation length from 83.5 to 139.8. By contrast, both BoN by Prob-Drop and BoN by Prob-Hold improve recall without sacrificing precision. Finally, by combining both Prob-Drop and Prob-Hold into our final SelfCite reward, we achieve the best outcome, increasing both recall and precision and a 4-point improvement in F1.
We also explored different token-length limits for citations in the bottom of Table 3, as discussed in Section 2.3. By default, we exclude candidates citing more than 384 tokens, unless the citation contains only a single sentence. Lowering the cap to 256 tokens slightly hurts F1, while raising it to 512 tokens has negligible impact. Completely removing length limits inflates citation length to 121.9 tokens and yields worse precision (79.3) but slightly improved recall (67.9). We also notice that the 256 length limit still outperforms the LongCite-8B baseline (66.4 vs 64.1) while having almost equally long citation length (84.5 vs 83.5), showing that the improvement of SelfCite correlates less with the citation length. Overall, using a 384-token limit achieves a good balance for short citation lengths and strong performance.
Table 3: Ablation study on HotpotQA citation recall, precision, and F1 (R, P, F1) and citation length for BoN decoding methods.
Decoding Methods HotpotQA Citation
R P F1 Length
LongCite-8B (Our repro.) 60.8 77.9 64.1 83.5
+ BoN by LM log prob 62.7 75.5 63.4 74.6
+ BoN by max citation length 66.5 73.6 65.1 139.8
+ BoN by Prob-Drop 65.6 78.1 66.6 92.9
+ BoN by Prob-Hold 66.2 78.1 67.0 93.4
+ BoN by SelfCite 67.2 81.0 68.8 93.4
w/ lower length limit (256) 65.8 78.8 66.4 84.5
w/ higher length limit (512) 67.0 82.2 68.5 99.2
w/o length limit ($\infty$) 67.9 79.3 68.1 121.9
Table 4: Ablation study on HotpotQA citation recall, precision, and F1 (R, P, F1) and citation length for finetuned models.
Fine-tuning Methods HotpotQA Citation
R P F1 Length
LongCite-8B (Our repro.) 60.8 77.9 64.1 83.5
+ SimPO 69.4 82.3 71.5 105.7
+ SimPO + BoN 72.0 82.7 72.9 126.9
+ SimPO w/ or w/o length balancing
w/ length balancing 69.4 82.3 71.5 105.7
w/o length balancing 64.4 62.9 60.5 152.9
+ SimPO w/ varying data sizes
1K examples 62.5 78.9 65.7 90.1
2K examples 69.4 82.3 71.5 105.7
4K examples 68.5 80.4 70.3 134.1
8K examples 63.6 75.3 64.4 195.2
+ SFT on BoN responses 68.8 77.3 68.4 98.7
+ SimPO by denoising perturbed citations
On original responses 40.5 50.5 41.6 88.8
On BoN responses 42.6 50.7 42.3 79.7
4.2 Citation Length Balance
As noted in Section 3.2, BoN selects slightly longer citations, making it easy for a model trained directly on BoN-preferred data to adopt the shortcut of generating longer citations without improving quality. To counter this, we apply length balancing, injecting random citations into examples where length bias exists to equalize the number of cited sentences. Table 4 (see w/ vs. w/o length balancing) highlights its critical role in length balancing. Without length balancing, the model overextends citations (average length 152.9), leading to lower precision (62.9) and F1 (60.5). In contrast, enabling length balancing maintains high precision (82.3) and recall (69.4), achieving a better F1 of 71.5 while keeping citation length reasonable (105.7). These results confirm that length balancing prevents shortcut learning, ensuring the model truly learns to cite accurately.
4.3 Training Size of SimPO
In prior study (Zhou et al., 2023), 1K examples are sufficient to align user preferences effectively. Table 4 presents SimPO results with 1K to 8K examples. 1K examples already bring a moderate improvement, raising F1 from 64.1 to 65.7, with gains in precision and recall. Using 2K examples further boosts F1 to 71.5, while 4K leads to saturated improvement. However, at 8K examples, performance declines, and citation length rises to 195.2. We attribute this to SimPO’s off-policy nature, especially because it lacks a reference model to constrain the output distributions to be similar to the collected data. As training steps grow, the model may drift from the collected data, potential overfitting to the biases in preference data. Thus, further fine-tuning may degrade citation quality. To address this, we show initial results from iterative SimPO in Section 4.6.
4.4 SimPO vs. SFT on Best-of-N responses
We also show the effect of applying standard supervised fine-tuning (SFT) on the responses selected by best-of-N sampling, which is a simplified alternative of preference optimization. As the result shown in the last row in Table 4, SFT also improves the F1 score from 64.1 to 68.4, but it still falls behind 71.5 of SimPO. This result confirms that it is necessary to train the model via SimPO with preference data, which enables the model to distinguish between bad and good citations, and thus improve the citation quality.
4.5 Off-policy Denoising Perturbed Citations
We explored a purely off-policy alternative approach. Specifically, given a model-generated response, we randomly shift its citation spans to create perturbed variants. SimPO training pairs were then constructed by preferring the original citation over the perturbed one, encouraging the model to “denoise” citations by restoring their original spans. However, as shown at the bottom of Table 4, this approach degrades performance, both when applied to original and best-of-N responses. We attribute this to a mismatch between the training data and the model’s natural error distribution—since random shifts do not reflect typical citation errors, they fail to provide useful guidance for improvement.
Figure 2: Iteratively applying SimPO for three iterations.
4.6 Iterative Preference Optimization
It has been discussed that an on-policy alignment process can be beneficial to avoid reward exploitation (Bai et al., 2022) and maintains consistency between the generated data and the model’s evolving output distribution. We thus experiment with iteratively performing SimPO, similar to the concepts of recent studies (Pang et al., 2024; Yasunaga et al., 2024), to maintain the consistency between the generated data and the model’s evolving output distribution. Specifically, after fine-tuning with SimPO, we generate a new dataset via BoN, which is also 2K in size but not overlapped with previous iterations. We continue training the model and repeat the process for three rounds. As shown in Figure 2, while the largest improvement occurs in the first round, improvements continue over three iterations, which further validates the reliability of our reward signal. Iterative SimPO is still not perfect since it remains an off-policy method. Given that our reward can be cheaply computed, we believe that on-policy methods like PPO (Schulman et al., 2017) could further enhance performance. We leave the exploration of such approaches for future work.
4.7 Qualitative Study
Finally, we examine an example that requires citing multiple context sentences to support a complex response. As shown in Table 5, the response integrates information from sentences 302, 303, and 306. Direct sampling (2) omits sentence 302 while incorrectly including 305. In contrast, the best-of-N candidate (1) correctly includes 302 and excludes 305, achieving a slightly higher reward (0.578 vs. 0.547), demonstrating the effectiveness of our reward design. We also present candidates (3) and (4), which cite more irrelevant sentences and miss key citations, leading to even lower rewards. Additional qualitative examples are provided in Appendix 11.
Table 5: An example of differences in the citation from baseline vs BoN. Related information are highlighted in the context/response.
Sent. ID Context Sentences (only showing a paragraph due to limited space)
302 (✓) In general, consumer advocates believe that any comprehensive federal privacy policy should complement, and not supplant, sector-specific privacy legislation or state-level legislation.
303 (✓) Finding a global consensus on how to balance open data flows and privacy protection may be key to maintaining trust in the digital environment and advancing international trade.
304 (✗) One study found that over 120 countries have laws related to personal data protection.
305 (✗) Divergent national privacy approaches raise the costs of doing business and make it harder for governments to collaborate and share data, whether for scientific research, defense, or law enforcement.
306 (✓) A system for global interoperability in a least trade-restrictive and nondiscriminatory way between different national systems could help minimize costs and allow entities in different jurisdictions with varying online privacy regimes to share data via cross-border data flows.
Query Please write a one-page summary of the above government report.
Response (only single statement due to space) […] The report concludes by noting that finding a global consensus on how to balance open data flows and privacy protection may be key to maintaining trust in the digital environment and advancing international trade. The report suggests that Congress may consider comprehensive privacy legislation and examine the potential challenges and implications of building a system of interoperability between different national privacy regimes. […]
BoN Candidates Citation Strings (green: correct; red: wrong) Missing Citations SelfCite Reward
(1) Best candidate [302-303][306-306] 0.578
(2) Direct sampling [303-303][305-306] (302) 0.547
(3) Other candidate [303-304][308-308][310-311] (302, 306) 0.461
(4) Other candidate [303-303][309-309][311-311] (302, 306) 0.375
5 Related Work
Citations for Language Models.
Recent work has explored various approaches to teaching language models to generate citations, including fine-tuning with direct human feedback or annotations (Nakano et al., 2021; Menick et al., 2022; Slobodkin et al., 2024), annotation/feedback from external models (Huang et al., 2024), and prompting-based methods (Gao et al., 2022, 2023b) to explicitly incorporate relevant retrieved documents. To avoid the human annotation processes, Zhang et al. (2024) introduced CoF (“Coarse to Fine”), an automated multi-stage pipeline that simulates human annotations. This approach leverages proprietary LLMs for chunk-level retrieval and sentence-level citation extraction, achieving high citation quality through supervised fine-tuning. However, it depends on two powerful proprietary APIs—GLM-4 for the LLM and Zhipu Embedding-v2 for retrieval [^2] https://open.bigmodel.cn/pricing—with carefully designed prompting, effectively distilling the capabilities of these powerful proprietary APIs into much smaller models in 8B/9B. In contrast, our SelfCite aims at completely eliminating the reliance on annotations for citation, either from human or proprietary APIs. Instead, our method enables a small 8B model to assess citation quality itself using self-supervised reward signal from context ablation, effectively self-improving without external supervision.
Contributive Context Attribution.
Besides being self-supervised, SelfCite also adopts the view that citations should reference the sources from the context that a model actually uses when generating a statement–known as contributive attribution (Worledge et al., 2023)–rather than any sources that merely support the claim. Our reward signal naturally aligns with this attribution framework, as context ablation identifies the sources that cause the model to produce a statement. Existing contributive attribution methods for LLMs typically require extensive context ablations or other computationally expensive techniques, such as gradient-based analysis during inference (Cohen-Wang et al., 2024; Qi et al., 2024; Phukan et al., 2024). In contrast, SelfCite simply generate the citation tags, and refine citation candidates by preference optimization with reward signals from context ablations, effectively teaching the model to perform contributive context attribution itself.
Self- or Weakly-Supervised Alignment.
Another relevant area is self- or weakly-supervised approaches for aligning LLMs without human supervision (Kim et al., 2023; Yuan et al., 2024), reducing the need for explicit human feedback (Ouyang et al., 2022), or curating high-quality data for supervised fine-tuning (Zhou et al., 2023). SelfCite shares the same spirit by computing simple probability differences under context ablation as rewards, eliminating the need for additional annotation process.
6 Conclusion and Limitations
In this work, we introduced SelfCite, a self-supervised framework that aligns LLMs to generate more accurate, fine-grained citations by directly leveraging their own probabilities for necessity and sufficiency rewards through context ablation. With best-of-N sampling and preference optimization, SelfCite significantly improves citation correctness on the LongBench-Cite benchmark without requiring human annotation, offering a promising self-improving direction towards verifiable and trustworthy LLMs.
SelfCite also has limitations: 1) While achieving good results with SimPO, integrating other alignment algorithms remains unexplored. 2) While we focus on self-supervision in preference optimization, our attempt to make the SFT stage self-supervised via ContextCite is still preliminary. Better unsupervised ways to kick-start LLMs’ ability in generating structured citations can be further explored.
Acknowledgements
We thank Jiajie Zhang and Yushi Bai for their help in providing implementation details of LongCite. We also thank Andrei Barbu, Linlu Qiu, Nour Jedidi, Weijia Shi, and Tianyu Gao for their valuable discussions. At MIT, Yung-Sung was sponsored by the United States Air Force Research Laboratory and the United States Air Force Artificial Intelligence Accelerator and was accomplished under Cooperative Agreement Number FA8750-19-2-1000. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the United States Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.